Suppose there are some low-quality firms and some high-quality firms which considers of issuing shares in initial public offering, and there are sufficient potential buyers who are considering of purchasing the shares. Each stock issuer values each lot of shares at $8,000 in the normal case, and each of them values each lot of shares at $1,000 if there will be a firm specific event after the share issuance. Each of the potential buyer values the shares at $9,000 per lot in the normal case, and $100 with a firm specific event. A low-quality firm will be at the normal case with probability of 0.6, and it would face a firm specific event with a probability of 0.4. A high-quality firm be at the normal case with probability of 0.9, and it would face a firm specific event with a probability of 0.1. It is assumed that everyone only cares the expected value and does not care the risk.

(a) Suppose only the stock issuers can observe the own quality of each frim, and each potential buyer only knows that the proportion of high quality firms in the economy is at 30% and the rest are low quality, would there be any adverse selection problem in this case, and would there be any initial public offering by each firm? If so, what is the range of the potential price if transaction costs are ignored? Calculation and explanation (with not more than 180 words) are required to acquire full credits. (8 marks)

(b) Suppose the government decides to force each firm to disclose a certain information, in which the potential stock purchasers can accurately expect the probability of having a firm specific event by each firm, would all firms participate in the initial public offering ? Calculation and explanation (with not more than 150 words) are required to acquire full credits. (6 marks)

(c) Explain (with not more than 150 words) the implications for the importance of a certain financial regulations based on your answer in part (a) and part (b). (6 marks)
Answer #1

(a) It is assumed that the probability of high quality firm is 30% and 70% will be low quality firm.

Now the adverse selection problem refers to the fact that the selection will be different of what it should have been. Therefore the price range will be as per stock issuer be

For low quality firm = 0.70*0.6*8000 = 3360

For high quality firm = 03*0.9*8000=2160.

Therefore the price range would be $2160 to $ 3360

(b) Now if the firm expects to force each firm to disclose certain information, the probability of certain events would be as follows:

For low quality firm = 0.7*0.4*1000 = 280

For high quality firm - 0.3*0.1*100 = 300

Therefore the range will be $280 to $300 in case of specific events.

(c) Financial regulation is important for finding the correct reported numbers by the firms. This helps in correct reporting by the company and also the true and clear picture of the economy could be devised. The financial implication in normal cases and specific cases help in making correct assumptions about the reported facts by the company.